# Reinforcement Learning via Self-Distillation

### v1

@@ 4.6 Which feedback is most informative?
$f=$ output $32.5\pm 0.5$ $13.7\pm 0.6$ $39.8\pm 0.2$ $0.40\pm 0.0$
$f=$ solution $\mathbf{42.4}\pm 1.0$ $12.1\pm 0.7$ $36.8\pm 2.7$ $\emph{0.07}\pm 0.0$
$f=$ output + solution $\mathbf{42.5}\pm 1.2$ $\mathbf{10.1}\pm 0.2$ $\mathbf{48.9}\pm 0.9$ $0.37\pm 0.0$
$f=$ $y$ + output + solution $39.3\pm 0.8$ $30.0\pm 0.9$ $44.5\pm 1.8$ $\emph{0.23}\pm 0.0$
Table 6: Performance of varying kinds of feedback. We evaluate informativeness of feedback based on SDPO training (until step 70) as well as the direct impact on the self-teacher. “Same output” measures the percentage of cases where the teacher receives the same environment output as the student’s initial attempt (i.e., not exploring alternative approaches). We observe that environment output and sample solutions are complementary and each provide informative feedback. Naively including only solutions or initial attempts $y$ significantly reduces diversity in the teacher and student. We remark that the sample solutions are generated by the student, enabling similar group-relative advantage estimation to GRPO. Error bars indicate standard deviation across 3 seeds.
@@ 6.1 Reinforcement Learning with LLMs
Conceptually, our work is related to “expert iteration” (Anthony et al., 2017) where a student is bootstrapped by repeatedly imitating an improved version of itself (called the “expert”). Canonically, the expert combines the student with test-time search, such as tree search (Anthony et al., 2017) or majority voting (Zuo et al., 2025). In contrast, SDPO leverages the student’s ability to learn from rich feedback provided in-context.
@@ 6.2 Learning from Rich Feedback and through Retrospection
Beyond scalar outcome rewards, recent works have leveraged rich execution or verbal feedback to guide generation (Gehring et al., 2025; Yuksekgonul et al., 2025). A primary line of research focuses on translating verbal feedback into reward functions for RL. This is often achieved by mapping feedback to discrete token-level rewards using an external frozen model (Wang et al., 2026), or by employing strong external LLMs to explicitly construct state-wise reward functions (Goyal et al., 2019; Xie et al., 2024; Urcelay et al., 2026).
@@ 6.3 Distillation
Distillation is frequently employed as an alternative to supervised fine-tuning (SFT) when a strong teacher model is available. This approach transfers capabilities by training a student to mimic the output distribution or intermediate representations of the teacher (Hinton et al., 2015; Romero et al., 2015; Kim & Rush, 2016; Sanh et al., 2019; Xie et al., 2020). Distillation is typically performed on fixed off-policy datasets. To address the distribution shift between training and inference, recent works explore on-policy distillation, where the student learns from feedback of an external teacher on its own generations (Agarwal et al., 2024; Gu et al., 2024; Yang et al., 2025; Lu & Thinking Machines Lab, 2025). This mitigates the train-test mismatch, which relates closely to earlier work on online imitation learning (Ross et al., 2011).
@@ 6.4 Self-Distillation
The concept of self-distillation was first proposed by Snell et al. (2022) in a setting akin to supervised learning, introducing the idea of sampling from a model provided with extra context and training the same model to mimic these predictions without that context. This mechanism has proven effective for compressing behavior (Bai et al., 2022; Choi et al., 2022) and factual information (Eyuboglu et al., 2026; Kujanpää et al., 2025) into model weights. Beyond compressing a fixed context into model weights, recent works have used self-distillation to learn from environment feedback (Scheurer et al., 2023; Dou et al., 2024; Mitra & Ulukus, 2025). These approaches use an off-policy self-distillation objective, which substantially underperforms SDPO’s on-policy learning. Off-policy self-distillation trains the student on generations from the teacher, whereas SDPO trains the student to avoid mistakes in its own generations. In concurrent work, Chen et al. (2025c) apply on-policy self-distillation to grid world settings where feedback is a scalar reward, and a reflection stage in the self-teacher diagnoses possible mistakes, showing improved credit assignment compared to learning value networks for advantage estimation.
@@ 7 Conclusion, Limitations, and Future Work
Empirically, SDPO demonstrates superior sample efficiency and wall-clock convergence compared to GRPO on reasoning tasks, even when training in standard RLVR environments without rich feedback. SDPO’s gains grow with model scale, suggesting that the capacity for self-correction scales with the model’s in-context learning capabilities. Moreover, we show that performing SDPO at test-time on individual hard binary-reward tasks accelerates the discovery of solutions compared to strong baselines.
SDPO enables learning from rich feedback in a way that parallels human cognition: utilizing precise outcomes rather than just binary rewards. By allowing the model to determine retrospectively how it should have acted, we demonstrate that language models can convert diverse tokenized feedback into effective self-supervision.
@@ Acknowledgments
We would like to thank Akira Yoshiyama, Yassir Akram, Parnian Kassraie, Jonathan Thomm, Roman Vorushin, Afra Amini, Imanol Schlag, Yu Sun, and Moritz Hardt for helpful discussions. We thank Eduard Durech for helpful conversations regarding the scaling of RL fine-tuning and for his technical guidance on distributed infrastructure and long-context optimization. Furthermore, we would like to thank Leander Diaz-Bone for supporting dataset generation.
@@ Appendix A Implementation of SDPO
Figure 14: The pseudo-code of SDPO within a standard RL training pipeline. Omitted here is the filtering to top-$K$ logprobs for student and teacher (including a tail term) as described in Section˜A.2. Further, we omit here any importance sampling weights to correct for off-policy data. reprompt modifies the batch to incorporate teacher context (i.e., rich feedback). divergence implements any per-token divergence such as reverse-KL, forward-KL, or Jensen-Shannon.
@@
  • Teacher regularization (Section˜A.1)
  • Approximating logit-distillation with the top-$K$ logits for saving GPU memory (Section˜A.2)
  • Generalizing PPO-style policy gradient algorithms to logit-level advantages (Section˜A.3)
@@
A.1 Regularized teacher
@@
    Exponential moving average (EMA): Alternatively, we can stabilize the teacher’s parameters directly; parameterizing $q_{\theta^{\prime}}$ by $\theta^{\prime}$ and updating as $\theta^{\prime}\leftarrow(1-\alpha)\theta^{\prime}+\alpha\theta$ with $\alpha\in(0,1)$. Under mild smoothness assumptions, this EMA teacher remains implicitly within a trust-region around the initial teacher (cf. Section˜B.3).
@@
A.2 Approximate Logit Distillation
@@
$\mathcal{L}_{\mathrm{SDPO}}(\theta)$ $=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}{}\sum_{t=1}^{T}\mathrm{KL}(\pi_{\theta}(y_{t}\mid x,y_{<t})\|\mathrm{stopgrad}(q_{\theta}(y_{t}\mid x,f,y_{<t})))$
@@
A.3 Off-Policy Training: Generalization to Logit-Level Losses
@@
where $\hat{y}_{i,t}$ sums over all possible tokens at position $t$ for rollout $i$ (or the $K$ most likely under $\pi_{\theta_{{\mathrm{old}}}}$, cf. Section˜A.2). The TIS changes since we explicitly weight each logit by its probability under $\pi_{\theta_{{\mathrm{old}}}}$ rather than relying on a Monte Carlo estimate of the expectation over next-token predictions. Here, $A_{i,t}(\hat{y}_{i,t})$ is a per-logit advantage.
@@ Appendix B Theoretical Analysis
B.1 Gradient Estimator
Proof of Proposition 2.1.
@@ B.1 Gradient Estimator
$\boldsymbol{\nabla}_{\!\!\theta}\,\mathcal{L}_{\mathrm{SDPO}}(\theta)$ $=\boldsymbol{\nabla}_{\!\!\theta}\,\sum_{t=1}^{T}\mathrm{KL}(\pi_{\theta}(y_{t}\mid x,y_{<t})\|\mathrm{stopgrad}(q_{\theta}(y_{t}\mid x,f,y_{<t})))$
$=\boldsymbol{\nabla}_{\!\!\theta}\,\sum_{t=1}^{T}\sum_{y_{t}}\pi_{\theta}(y_{t}\mid x,y_{<t})\log\left(\frac{\pi_{\theta}(y_{t}\mid x,y_{<t})}{\mathrm{stopgrad}(q_{\theta}(y_{t}\mid x,f,y_{<t}))}\right)$
Let $A_{t,k}:=\log\left(\frac{\mathrm{stopgrad}(q_{\theta}(y_{t}\mid x,f,y_{<t}))}{\pi_{\theta}(y_{t}\mid x,y_{<t})}\right)$. Then,
$=-\boldsymbol{\nabla}_{\!\!\theta}\,\sum_{t=1}^{T}\sum_{y_{t}}\pi_{\theta}(y_{t}\mid x,y_{<t})A_{t,k}$
$=-\sum_{t=1}^{T}\sum_{y_{t}}\pi_{\theta}(y_{t}\mid x,y_{<t})\boldsymbol{\nabla}_{\!\!\theta}\,A_{t,k}+A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(y_{t}\mid x,y_{<t}).$
We have that $\boldsymbol{\nabla}_{\!\!\theta}\,A_{t,k}=-\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(y_{t}\mid x,y_{<t})$ is the negative score function. Using the score trick, $\pi_{\theta}(y_{t}\mid x,y_{<t})\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(y_{t}\mid x,y_{<t})=\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(y_{t}\mid x,y_{<t})$. Hence, the first term simplifies to
$-\sum_{t=1}^{T}\sum_{y_{t}}\pi_{\theta}(y_{t}\mid x,y_{<t})\boldsymbol{\nabla}_{\!\!\theta}\,A_{t,k}$ $=\sum_{t=1}^{T}\sum_{y_{t}}\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(y_{t}\mid x,y_{<t})=\sum_{t=1}^{T}\boldsymbol{\nabla}_{\!\!\theta}\,\underbrace{\sum_{y_{t}}\pi_{\theta}(y_{t}\mid x,y_{<t})}_{=1}=0.$
@@
$\boldsymbol{\nabla}_{\!\!\theta}\,\mathcal{L}_{\mathrm{SDPO}}$ $=-\sum_{t=1}^{T}\sum_{y_{t}}A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(y_{t}\mid x,y_{<t})$
$=-\sum_{t=1}^{T}\sum_{y_{t}}\pi_{\theta}(y_{t}\mid x,y_{<t})\Big(A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(y_{t}\mid x,y_{<t})\Big)$
$=-\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\sum_{t=1}^{|y|}\sum_{y_{t}}\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(y_{t}\mid x,y_{<t})\cdot A_{t,k}\right]$
$=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\sum_{t=1}^{|y|}\sum_{y_{t}}\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(y_{t}\mid x,y_{<t})\cdot\log\frac{\pi_{\theta}(y_{t}\mid x,y_{<t})}{q_{\theta}(y_{t}\mid x,f,y_{<t})}\right].$
Notably, the above implies that the gradient of $\mathcal{L}_{\mathrm{SDPO}}$ is equivalent to the gradient of the loss if $A_{t,k}=\mathrm{stopgrad}\left(\log\frac{\pi_{\theta}(y_{t}\mid x,y_{<t})}{q_{\theta}(y_{t}\mid x,f,y_{<t})}\right)$.

### v2

@@ 4.6 Which feedback is most informative?
$f=$ output $32.5\pm 0.5$ $13.7\pm 0.6$ $39.9\pm 1.1$ $0.40\pm 0.0$
$f=$ own solution $\mathbf{42.4}\pm 1.0$ $12.1\pm 0.7$ $42.6\pm 1.3$ $0.41\pm 0.0$
$f=$ output + own solution $\mathbf{42.5}\pm 1.2$ $\mathbf{10.1}\pm 0.2$ $\mathbf{48.3}\pm 1.4$ $0.38\pm 0.0$
$f=$ $y$ + output + own solution $39.3\pm 0.8$ $30.0\pm 0.9$ $44.5\pm 1.3$ $\emph{0.23}\pm 0.0$
Table 6: Performance of varying kinds of feedback. We evaluate informativeness of feedback based on SDPO training (until step 60) as well as the direct impact on the self-teacher. “Same output” measures the percentage of cases where the teacher receives the same environment output as the student’s initial attempt (i.e., not exploring alternative approaches). We observe that environment output and sample solutions are complementary and each provide informative feedback. Naively including only solutions or initial attempts $y$ significantly reduces diversity in the teacher and student. We remark that the sample solutions are generated by the student, enabling similar group-relative advantage estimation to GRPO. Error bars indicate standard deviation across 3 seeds.
@@ 6.1 Reinforcement Learning with LLMs
Conceptually, our work is related to “bootstrapping your own latent” (BYOL; Grill et al., 2020) and “expert iteration” (Anthony et al., 2017) where a student is bootstrapped by repeatedly imitating an improved version of itself (called the “expert”). Canonically, the expert combines the student with test-time search, such as tree search (Anthony et al., 2017) or majority voting (Zuo et al., 2025). In contrast, SDPO leverages the student’s ability to learn from rich feedback provided in-context, which is related to “augmented views” in BYOL.
@@ 6.2 Learning from Rich Feedback and through Retrospection
Beyond scalar outcome rewards, recent works have leveraged rich execution or verbal feedback to guide generation (Gehring et al., 2025; Feng et al., 2024b; Yuksekgonul et al., 2025). A primary line of research focuses on translating verbal feedback into reward functions for RL. This is often achieved by mapping feedback to discrete token-level rewards using an external frozen model (Wang et al., 2026), or by employing strong external LLMs to explicitly construct state-wise reward functions (Goyal et al., 2019; Xie et al., 2024; Urcelay et al., 2026).
@@ 6.3 Distillation
Distillation is frequently employed as an alternative to supervised fine-tuning (SFT) when a strong teacher model is available. Distillation transfers capabilities by training a student to mimic the output distribution or intermediate representations of the teacher (Hinton et al., 2015; Romero et al., 2015; Kim & Rush, 2016; Sanh et al., 2019; Xie et al., 2020). While often performed on fixed off-policy datasets, to address the distribution shift between training and inference, recent works explore on-policy distillation, where the student learns from feedback on its own generations provided by an external teacher (Agarwal et al., 2024; Gu et al., 2024; Yang et al., 2025a; Lu & Thinking Machines Lab, 2025). This mitigates the train-test mismatch, which relates closely to earlier work on online imitation learning (Ross et al., 2011).
@@ 6.4 Self-Distillation
The concept of self-distillation was first proposed by Snell et al. (2022) in a setting akin to supervised learning, introducing the idea of sampling from a model provided with extra context and training the same model to mimic these predictions without that context. This mechanism has proven effective for compressing behavior (Bai et al., 2022; Choi et al., 2022; Yang et al., 2024; 2025b) and factual information (Eyuboglu et al., 2026; Kujanpää et al., 2025; Cao et al., 2025a) into model weights. Beyond compressing a fixed context into model weights, recent works have used self-distillation to learn from environment feedback (Scheurer et al., 2023; Dou et al., 2024; Zhou et al., 2025; Mitra & Ulukus, 2025; Song et al., 2026). These approaches use an off-policy self-distillation objective, which we find to substantially underperform SDPO’s on-policy learning. Off-policy self-distillation trains the student on generations from the teacher, whereas SDPO trains the student to avoid mistakes in its own generations. In concurrent work, Chen et al. (2025c) apply on-policy self-distillation to grid world settings where feedback is a scalar reward, and a reflection stage in the self-teacher diagnoses possible mistakes, showing improved credit assignment compared to learning value networks for advantage estimation. Other concurrent work studies SDPO on a fixed dataset of expert demonstrations, without online environment interaction (Shenfeld et al., 2026a; Zhao et al., 2026).
@@ 7 Conclusion, Limitations, and Future Work
Empirically, SDPO demonstrates superior sample efficiency and wall-clock convergence compared to GRPO on reasoning tasks, even when training in standard RLVR environments without rich feedback. SDPO’s gains grow with model scale, suggesting that the capacity for self-correction scales with the model’s in-context learning capabilities. Moreover, we show that performing SDPO at test time on individual hard binary-reward tasks accelerates the discovery of solutions compared to strong baselines.
SDPO enables learning from rich feedback in a way that is arguably closer to human cognition: utilizing precise outcomes rather than just binary rewards. By allowing the model to determine retrospectively how it should have acted, we demonstrate that language models can convert diverse tokenized feedback into effective self-supervision.
@@ Acknowledgments
We would like to thank Akira Yoshiyama, Yassir Akram, Parnian Kassraie, Jonathan Thomm, Roman Vorushin, Afra Amini, Imanol Schlag, Yu Sun, and Moritz Hardt for helpful discussions. We thank Eduard Durech for helpful conversations regarding the scaling of RL fine-tuning and for his technical guidance on distributed infrastructure and long-context optimization. We are grateful to Ruixu Zhou from Tsinghua University & the Tencent Hunyuan Team for pointing out an error in the initially derived gradient estimator. Furthermore, we would like to thank Leander Diaz-Bone for supporting dataset generation.
@@ Appendix A Implementation of SDPO
Figure 14: The pseudo-code of SDPO within a standard RL training pipeline. Omitted here is the filtering to top-$K$ logprobs for student and teacher (including a tail term) as described in Section˜A.3. Further, we omit here any importance sampling weights to correct for off-policy data. reprompt modifies the batch to incorporate teacher context (i.e., rich feedback). divergence implements any per-token divergence such as reverse-KL, forward-KL, or Jensen-Shannon.
@@
  • The gradient estimator used in our implementation (Section˜A.1)
  • Teacher regularization (Section˜A.2)
  • Approximating logit-distillation with the top-$K$ logits for saving GPU memory (Section˜A.3)
  • Generalizing PPO-style policy gradient algorithms to logit-level advantages (Section˜A.4)
@@
A.1 Gradient Estimators
In this seciton, we discuss two possible gradient estimators for the KL divergence between the current policy $\pi_{\theta}(y\mid x)$ and the teacher policy $q_{\theta}(y\mid x,f)$.
Per-token estimator.
Deriving the gradient of the SDPO loss as defined in Equation˜1:
$\mathcal{L}_{\mathrm{token}}(\theta):=\mathbb{E}_{y\sim\mathrm{stopgrad}(\pi_{\theta}(\cdot\mid x))}\left[\sum_{t=1}^{T}\mathrm{KL}(\pi_{\theta}(\cdot\mid x,y_{<t})\|\mathrm{stopgrad}(\pi_{\theta}(\cdot\mid x,f,y_{<t})))\right]$ (5)
leads to the following estimator (see a detailed proof in Section˜B.1), which corresponds to the sum of gradients of the KL divergence at each token:
$\boldsymbol{\nabla}\mathcal{L}_{\text{token}}(\theta)=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\sum_{t=1}^{T}\mathbb{E}_{\hat{y}_{t}\sim\pi_{\theta}(\cdot\mid x,y_{<t})}\!\!\left[\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\cdot\log\frac{\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})}{\pi_{\theta}(\hat{y}_{t}\mid x,f,y_{<t})}\right]\right].$ (6)
This corresponds to the estimator presented in Proposition 2.1. This gradient estimator effectively assumes that the sampling distribution generating $y$ is fixed.
Sequence-level estimator.
An alternative self-distillation objective minimizes the sequence-level KL divergence between student and self-teacher, i.e.,
$\begin{split}\mathcal{L}_{\mathrm{seq}}(\theta):=\mathrm{KL}\left(\pi_{\theta}\|q_{\theta}\right)&=\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\log\frac{\pi_{\theta}(y\mid x)}{q_{\theta}(y\mid x,f)}\right]\\ &=\sum_{t=1}^{T}\mathbb{E}_{s_{t}\sim\Pi_{\theta}}\left[\mathrm{KL}\left(\pi_{\theta}(\cdot\mid s_{t})\|q_{\theta}(\cdot\mid s_{t},f)\right)\right],\end{split}$ (7)
where $s_{t}=(x,y_{<t})$ is the prefix (“state”) at step $t$ and $\Pi_{\theta}$ denotes the prefix distribution under policy $\pi_{\theta}$. Estimating the gradient of this objective additionally takes into account how the choice of $y_{t}$ influences future states $y_{>t}$ (due to the additional dependence on $\Pi_{\theta}$).
Amini et al. (2025) show that the corresponding gradient estimator is given by
$\boldsymbol{\nabla}\mathcal{L}_{\text{seq}}(\theta)=\boldsymbol{\nabla}\mathcal{L}_{\text{token}}(\theta)+\mathbb{E}_{y\sim\pi_{\theta}(\cdot\mid x)}\left[\sum_{t=1}^{T}\mathrm{KL}\left(\pi_{\theta}(\cdot\mid s_{t})\|q_{\theta}(\cdot\mid s_{t},f)\right)\boldsymbol{\nabla}_{\!\!\theta}\,\log\Pi_{\theta}(s_{t})\right].$ (8)
The additional term of the sequence-level gradient captures how prefixes influence the self-distillation divergence of future tokens. We also experimented with this sequence-level gradient estimator but did not find measurable gains relative to its additional complexity.
A.2 Regularized teacher
@@
    Exponential moving average (EMA): Alternatively, we can stabilize the teacher’s parameters directly; parameterizing $q_{\theta^{\prime}}$ by $\theta^{\prime}$ and updating as $\theta^{\prime}\leftarrow(1-\alpha)\theta^{\prime}+\alpha\theta$ with $\alpha\in(0,1)$.
@@
A.3 Approximate Logit Distillation
@@
$\mathcal{L}_{\mathrm{SDPO}}(\theta)$ $=\sum_{t=1}^{T}\mathrm{KL}(\pi_{\theta}(\cdot\mid x,y_{<t})\|\mathrm{stopgrad}(q_{\theta}(\cdot\mid x,f,y_{<t})))$
@@
A.4 Off-Policy Training: Generalization to Logit-Level Losses
@@
where $\hat{y}_{i,t}$ sums over all possible tokens at position $t$ for rollout $i$ (or the $K$ most likely under $\pi_{\theta_{{\mathrm{old}}}}$, cf. Section˜A.3). The TIS changes since we explicitly weight each logit by its probability under $\pi_{\theta_{{\mathrm{old}}}}$ rather than relying on a Monte Carlo estimate of the expectation over next-token predictions. Here, $A_{i,t}(\hat{y}_{i,t})$ is a per-logit advantage.
@@ Appendix B Theoretical Analysis
B.1 Proof of Proposition 2.1.
Proof.
@@
$\boldsymbol{\nabla}_{\!\!\theta}\,\mathcal{L}_{\mathrm{SDPO}}(\theta)$ $=\boldsymbol{\nabla}_{\!\!\theta}\,\sum_{t=1}^{T}\mathrm{KL}(\pi_{\theta}(\cdot\mid x,y_{<t})\|\mathrm{stopgrad}(q_{\theta}(\cdot\mid x,f,y_{<t})))$
$=\boldsymbol{\nabla}_{\!\!\theta}\,\sum_{t=1}^{T}\sum_{\hat{y}_{t}}\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\log\left(\frac{\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})}{\mathrm{stopgrad}(q_{\theta}(\hat{y}_{t}\mid x,f,y_{<t}))}\right)$
Let $A_{t,k}:=\log\left(\frac{\mathrm{stopgrad}(q_{\theta}(\hat{y}_{t}\mid x,f,y_{<t}))}{\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})}\right)$. Then,
$=-\boldsymbol{\nabla}_{\!\!\theta}\,\sum_{t=1}^{T}\sum_{\hat{y}_{t}}\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})A_{t,k}$
$=-\sum_{t=1}^{T}\sum_{\hat{y}_{t}}\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\boldsymbol{\nabla}_{\!\!\theta}\,A_{t,k}+A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t}).$
We have that $\boldsymbol{\nabla}_{\!\!\theta}\,A_{t,k}=-\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})$ is the negative score function. Using the score trick, $\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})=\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})$. Hence, the first term simplifies to
$-\sum_{t=1}^{T}\sum_{\hat{y}_{t}}\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\boldsymbol{\nabla}_{\!\!\theta}\,A_{t,k}$ $=\sum_{t=1}^{T}\sum_{\hat{y}_{t}}\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})=\sum_{t=1}^{T}\boldsymbol{\nabla}_{\!\!\theta}\,\underbrace{\sum_{\hat{y}_{t}}\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})}_{=1}=0.$
@@
$\boldsymbol{\nabla}_{\!\!\theta}\,\mathcal{L}_{\mathrm{SDPO}}$ $=-\sum_{t=1}^{T}\sum_{\hat{y}_{t}}A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})$
$=-\sum_{t=1}^{T}\sum_{\hat{y}_{t}}\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\Big(A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\Big)$
$=-\sum_{t=1}^{T}\mathbb{E}_{\hat{y}_{t}\sim\pi_{\theta}(\cdot\mid x,y_{<t})}\left[A_{t,k}\boldsymbol{\nabla}_{\!\!\theta}\,\log\pi_{\theta}(\hat{y}_{t}\mid x,y_{<t})\right].$
Notably, the above implies that the gradient of $\mathcal{L}_{\mathrm{SDPO}}$ is equivalent to the gradient of the loss if $A_{t,k}=\mathrm{stopgrad}\left(\log\frac{q_{\theta}(y_{t}\mid x,f,y_{<t})}{\pi_{\theta}(y_{t}\mid x,y_{<t})}\right)$.